OpenStudy (anonymous):

what is the sqr root of 1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

how do you no

OpenStudy (anonymous):

2

OpenStudy (anonymous):

1^2 = 1* 1 = 1

OpenStudy (anonymous):

you input into a calculator.